OpenStudy (anonymous):

How can I find the factors of (2x^2-9x-5). Please show steps.

OpenStudy (ranga):

Try the "AC" method. That is, in ax^2 + bx + c, multiply a*c split b into two parts so that it adds up to b and the product is ac. Here you have 2x^2 - 9x - 5 2 * (-5) = -10 Split the middle -9 so that the sum is -9 and the product is -10 -9 = -10 + 1. The sum is -9 and the product is -10 So 2x^2 - 9x - 5 = 2x^2 -10x + 1x - 5 Factor the first two terms and the last two terms: 2x(x - 5) + 1(x - 5) (x-5)(2x+1)

OpenStudy (anonymous):

I see, that makes sense. Thank you again ranga!

OpenStudy (ranga):

You are welcome.